OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function y = -3x2 + 12x – 7

OpenStudy (anonymous):

y'= -6 x + 12

OpenStudy (anonymous):

y'=0 when x=2

OpenStudy (anonymous):

(2, y(2)) is vertex

OpenStudy (anonymous):

(2,5) is the vertex

OpenStudy (anonymous):

If you do not know derivatives, you complete the square and obtain \[ y=-3 x^2+12 x-7=-3 \left(x^2-4 x+4-4\right)-7=-3 (x-2)^2+5 \] If x =2, y=5